High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor- hours. At the beginning of the year, the company provided the following estimates Department Direct labor-hours Machine-hours Fixed manufacturing overhead cost Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour Molding Painting 51,300 31,000 $184, 800 $477,090 32,500 88,000 $2.80 $4.80 Job 205 was started on August1 and completed on August 10. The companys cost records show the following information concerning the job Department MoldingPainting 79 340 Direct labor-hours Machine-hours Direct materials Direct labor cost 127 70 $ 942 1,140 $ 740 990 Required 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 21 units, what would be the unit product cost?
Answer #1

1) Departmental overhead rate :

Molding department = (184800/88000)+2.80 = $4.90 per machine hour

Painting department = (477090/51300)+4.80 = $14.10 per direct labor hour

2) Total overhead cost applied = (340*4.90+127*14.10) = $3456.70 or 3457

3) Total manufacturing cost = 942+1140+740+990+3456.70 = $7268.70

4) Unit product cost = 7268.70/21 = $346.13 per unit